# G-RAG: Knowledge Expansion in Material Science

### v1

@@ Abstract
In the field of Material Science, effective information retrieval systems are essential for facilitating research. Traditional Retrieval-Augmented Generation (RAG) approaches in Large Language Models (LLMs) often encounter challenges such as outdated information, hallucinations, limited interpretability due to context constraints, and inaccurate retrieval. To address these issues, Graph RAG integrates graph databases to enhance the retrieval process. Our proposed method processes Material Science documents by extracting key entities (referred to as MatIDs) from sentences, which are then utilized to query external Wikipedia knowledge bases (KBs) for additional relevant information. We implement an agent-based parsing technique to achieve a more detailed representation of the documents. Our improved version of Graph RAG called G-RAG further leverages a graph database to capture relationships between these entities, improving both retrieval accuracy and contextual understanding. This enhanced approach demonstrates significant improvements in performance for domains that require precise information retrieval, such as Material Science.
@@ 2.1 Graph RAG vs G-RAG
Graph RAG effectively merges the strengths of retrieval-based and generative methods to enhance LLMs’ capability to generate accurate, relevant, and contextually enriched responses peng2024graph . While supplying an LLM with text chunks from extensive documents may result in issues with context, factual precision, and language coherence, Graph RAG addresses these limitations by utilizing a knowledge graph as a source of structured, factual information edge2024local . The knowledge graph provides detailed entity information, including attributes and relationships, allowing the LLM to gain a deeper understanding and produce more informed, precise responses. In our G-RAG system, entity linking is a fundamental component, enabling the extraction of specific entities (key terms or concepts) from the text using an entity extractor like a Span Parser. These identified entities are then used to query an external retriever, which fetches relevant MatIDs and their corresponding information from a Wikipedia knowledge base [^2] https://huggingface.co/relik-ie/relik-reader-deberta-v3-large-re-wikipedia. This targeted retrieval process ensures that the selected MatIDs are highly relevant and accurate, thereby preserving the integrity of the constructed knowledge graph [^3] https://neo4j.com/developer-blog/entity-linking-relationship-extraction-relik-llamaindex/. Following this, an LLM formulates a query that is sent to the graph database. The graph database retrieves relevant information, which is processed by the LLM to generate a final, comprehensive response. The complete architecture of our G-RAG system is illustrated in Figure 1.
@@ 5 Conclusion and Future Work
Our findings indicate that integrating graph-based techniques and ensuring robust entity linking with external databases can significantly enhance the performance of the Graph RAG pipeline, particularly in terms of response relevance and accuracy. This approach also mitigates the challenge of maintaining relevance observed in standard Graph RAG implementations. Future work could include developing a larger knowledge base tailored to material science as an extended information source, as well as creating a material science-specific entity linking model. Additionally, establishing a comprehensive evaluation metric for Graph RAG would provide deeper insights into the process and its effectiveness. The code will be available at https://github.com/RadeenXALNW/G-RAG_1.0.
@@ Appendix A Appendix
A.1 Documents Parsing Method
@@
A.2 Entity Linking and Relation Extraction
@@
A.3 Examples from Our Dataset
@@ A.4.4 Context Relevancy
A.4.5 Analysis of Variance (ANOVA)
@@
A.5 Examples of Different Metrics

### v2

@@ Abstract
In the field of Material Science, effective information retrieval systems are essential for facilitating research. Traditional Retrieval-Augmented Generation (RAG) approaches in Large Language Models (LLMs) often encounter challenges such as outdated information, hallucinations, limited interpretability due to context constraints, and inaccurate retrieval. To address these issues, Graph RAG integrates graph databases to enhance the retrieval process. Our proposed method processes Material Science documents by extracting key entities (referred to as MatIDs) from sentences, which are then utilized to query external Wikipedia knowledge bases (KBs) for additional relevant information. We implement an agent-based parsing technique to achieve a more detailed representation of the documents. Our improved version of Graph RAG called G-RAG further leverages a graph database to capture relationships between these entities, improving both retrieval accuracy and contextual understanding. This enhanced approach demonstrates significant improvements in performance for domains that require precise information retrieval, such as Material Science. The code is available at https://github.com/RadeenXALNW/G-RAG_1.0.
@@ 2.1 Graph RAG vs G-RAG
Graph RAG effectively merges the strengths of retrieval-based and generative methods to enhance LLMs’ capability to generate accurate, relevant, and contextually enriched responses peng2024graph . While supplying an LLM with text chunks from extensive documents may result in issues with context, factual precision, and language coherence, Graph RAG addresses these limitations by utilizing a knowledge graph as a source of structured, factual information edge2024local . The knowledge graph provides detailed entity information, including attributes and relationships, allowing the LLM to gain a deeper understanding and produce more informed, precise responses. In our G-RAG system, entity linking is a fundamental component, enabling the extraction of specific entities (key terms or concepts) from the text using an entity extractor like a Span Parser. These identified entities are then used to query an external retriever, which fetches relevant MatIDs and their corresponding information from a Wikipedia knowledge base [^2] https://huggingface.co/relik-ie/relik-reader-deberta-v3-large-re-wikipedia. This targeted retrieval process ensures that the selected MatIDs are highly relevant and accurate, thereby preserving the integrity and relevancy of the constructed knowledge graph [^3] https://neo4j.com/developer-blog/entity-linking-relationship-extraction-relik-llamaindex/. Following this, an LLM formulates a query that is sent to the graph database. The graph database retrieves relevant information, which is processed by the LLM to generate a final, comprehensive response. We set a limit on the number of nodes retrieved to ensure the data fits within the context length of the LLM model A.1. The complete architecture of our G-RAG system is illustrated in Figure 1.
@@ 5 Conclusion and Future Work
Our findings indicate that integrating graph-based techniques and ensuring robust entity linking with external databases can significantly enhance the performance of the Graph RAG pipeline, particularly in terms of response relevance and accuracy. This approach also mitigates the challenge of maintaining relevance observed in standard Graph RAG implementations. Future work could include developing a larger knowledge base tailored to material science as an extended information source, as well as creating a material science-specific entity linking model. We also aim to explore this method in other domains where retrieving accurate, precise, and relevant information is essential. Additionally, establishing a comprehensive evaluation metric for Graph RAG would provide deeper insights into the process and its effectiveness.
@@ Appendix A Appendix
A.1 Node Selection Process for retrieving information
Let:
  • $q$ be the user’s question.
  • $D=(N,R)$ be the graph database, where $N$ is the set of nodes and $R$ is the set of relationships.
  • LLM be the Large Language Model used to generate the answer.
  • $\text{ContextLength}_{\text{max}}$ be the maximum context length allowed by the LLM.
After that, now let’s define the keyword extraction function $K:Q\rightarrow 2^{\Sigma^{*}}$:
$K(q)=\{\text{lemma}(t)\mid t\in\text{Tokens}(q),\ \text{POS}(t)\in S,\ |\text{lemma}(t)|>l_{\text{min}}\}$
where:
  • $\text{Tokens}(q)$ is the set of tokens from question $q$.
  • $\text{lemma}(t)$ is the lemmatized form of token $t$.
  • $\text{POS}(t)$ returns the part-of-speech tag of token $t$.
  • $S=\{\text{NOUN},\text{PROPN},\text{ADJ},\text{VERB}\}$.
  • $l_{\text{min}}$ is the minimum length of a lemma.
Here, the relevance function is $\text{Rel}:\Sigma^{*}\times\Sigma^{*}\rightarrow\{0,1\}$:
$\text{Rel}(s,k)=\begin{cases}1,&\text{if }k\subseteq\text{lower}(s)\\ 0,&\text{otherwise}\end{cases}$
where $\text{lower}(s)$ converts string $s$ to lowercase.
Now, let’s define the node selection function $N_{K}:2^{\Sigma^{*}}\times N\rightarrow 2^{N}$:
$N_{K}(K,N)=\left\{n\in N\mid\exists k\in K,\ \text{Rel}(n.\text{text},k)=1\right\}$
And, then comes the relationship selection function $R_{K}:2^{\Sigma^{*}}\times R\rightarrow 2^{R}$:
$R_{K}(K,R)=\left\{r\in R\mid\exists k\in K,\ \text{Rel}(r.\text{text},k)=1\right\}$
Apply limits to ensure the context fits within the LLM’s context length:
$|N_{K}|\leq N_{\text{max}},\quad|R_{K}|\leq R_{\text{max}}$
where $N_{\text{max}}$ and $R_{\text{max}}$ are determined based on $\text{ContextLength}_{\text{max}}$.
Here, the context construction function will be $C:2^{N}\times 2^{R}\rightarrow\Sigma^{*}$:
$C(N_{K},R_{K})=\text{Concat}\left(\left\{n.\text{text}\mid n\in N_{K}\right\}\cup\left\{r.\text{text}\mid r\in R_{K}\right\}\right)$
where Concat concatenates the text attributes into a single string.
After that, we can generate data using prompt generation.
Define the prompt generation function $P:Q\times\Sigma^{*}\rightarrow\Sigma^{*}$:
$P(q,C)=\text{Template}(q,C)$
where $\text{Template}(q,C)$ is a predefined template that incorporates the question $q$ and context $C$.
The final answer $a$ is obtained by passing the prompt $P(q,C)$ to the LLM:
$a=\text{LLM}(P(q,C))$
A.2 Documents Parsing Method
@@
A.3 Entity Linking and Relation Extraction
@@
A.4 Examples from Our Dataset
@@ A.5.4 Context Relevancy
A.5.5 Analysis of Variance (ANOVA)
@@
A.6 Examples of Different Metrics